# Learning Chemical Reaction Representation with Reactant-Product Alignment

### v1

@@ Abstract.
Organic synthesis stands as a cornerstone of chemical industry. The development of robust machine learning models to support tasks associated with organic reactions is of significant interest. However, current methods rely on hand-crafted features or direct adaptations of model architectures from other domains, which lacks feasibility as data scales increase or overlook the rich chemical information inherent in reactions. To address these issues, this paper introduces RAlign, a novel chemical reaction representation learning model tailored for a variety of organic-reaction-related tasks. By integrating atomic correspondence between reactants and products, our model discerns the molecular transformations that occur during the reaction, thereby enhancing the comprehension of the reaction mechanism. We have designed an adapter structure to incorporate reaction conditions into the chemical reaction representation, allowing the model to handle diverse reaction conditions and adapt to various datasets and downstream tasks, e.g., reaction performance prediction. Additionally, we introduce a reaction-center aware attention mechanism that enables the model to concentrate on key functional groups, thereby generating potent representations for chemical reactions. Our model has been evaluated on a range of downstream tasks, including reaction condition prediction, reaction yield prediction, and reaction selectivity prediction. Experimental results indicate that our model markedly outperforms existing chemical reaction representation learning architectures across all tasks. Notably, our model significantly outperforms all the baselines with up to 25% (top-1) and 16% (top-10) increased accuracy over the strongest baseline on USPTO_CONDITION dataset for reaction condition prediction. We plan to open-source the code contingent upon the acceptance of the paper.
1. Introduction
Organic synthesis has long been an essential component of the organic chemical industry, particularly within the pharmaceutical sector. Despite advancements in chemical synthesis technology, a range of tasks related to organic reactions remain challenging for humans, such as retrosynthesis planning and reaction condition recommendation. With the growth in computing power, data availability, and AI techniques, various models have been developed for organic chemistry, including graph-based (Rong et al., 2020; Li et al., 2021) and sequence-based (Irwin et al., 2022) models. However, current research mainly concentrates on molecular representation learning, and there is still a deficiency in effective reaction representation to tackle the relation of molecules during the complex reaction process. This work focuses on the backbone design for chemical reaction representation learning, with the aim of improving this situation.
@@ 1. Introduction
Existing chemical reaction representation learning methods can be roughly classified into two groups: fingerprint-based and deep-learning based. fingerprint-based methods use hand-craft fingerprints as molecular representations. These methods (Probst et al., 2022; Sandfort et al., 2020) employ various strategies to integrate the molecular representations of different components of a chemical reaction to derive a comprehensive chemical reaction representation. The representations are usually combined with conventional machine learning models, such as XGBoost (Chen and Guestrin, 2016), to tackle downstream tasks. While these approaches do not necessitate extensive computational resources and have proven effective with limited datasets, they may encounter performance limitations when scaling to larger datasets and more complex scenarios. This is attributed to the oversimplification of information inherent in the manually designed, statistically based features.
@@
Working towards more powerful chemical reaction representations, researchers have increasingly turned to deep learning techniques. Leveraging SMILES (Weininger, 1988), chemical reactions can be encoded into a string format, allowing the application of natural language processing methodologies to address the challenge of chemical reaction representation learning (Lu and Zhang, 2022; Schwaller et al., 2021b). Some studies (Maser et al., 2021; Han et al., 2024) employed graph neural networks for this task, capitalizing on the natural graph structure of molecules. However, most of these methods tend to directly apply existing frameworks from other domains or perform independent feature extraction for each component of the reaction followed by a simple aggregation, potentially overlooking the rich information in complex chemical reactions.
@@
Beyond the insufficient utilization of reaction information, current methods simply regard reagents as a component of reactants for encoding when integrating reaction conditions into the reaction representation. However, this approach does not allow the model to consider any non-molecular reaction conditions, including temperature and other environmental factors (Goodman, 2009). Furthermore, this narrow focus also impedes the applicability of current methods to datasets that provide experimental operations (e.g., stir and filter) in the form of natural language (Kearnes et al., 2021). Hence, we aim to develop a chemical reaction representation learning model that integrates a richer set of chemical information and is adaptable to various modalities of reaction conditions.
@@
To address the aforementioned shortcomings, we proposed RAlign, a powerful chemical reaction representation learning model for multiple downstream tasks. Reaction centers and the reaction process play pivotal roles in determining the outcome of the reaction (Schwaller et al., 2021b). Drawing inspiration from the imaginary transition structures of organic reactions (Fujita, 1986), we incorporate information fusion operations for corresponding atom pairs in reactants and products within our encoder. This approach explicitly models the chemical bond changes during reactions. Furthermore, we have proposed a reaction-center-aware decoder to assist the model in focusing on key functional groups. To accommodate various modalities of reaction conditions, we have employed an adapter structure to integrate these conditions into the chemical reaction representations. We have evaluated our model on a range of tasks, including reaction condition prediction, reaction yield prediction, and reaction selectivity prediction. Our model has achieved remarkable performance across all tasks, even surpassing the baselines with extensive pretraining. The contribution of this work can be summarized as:
@@
To the best of our knowledge, this work is the first to model atomic correspondence between reactants and products in the extraction of reaction representations, and it is also the first to design a graph backbone specifically for chemical reaction representation learning.
@@
Extensive experiments demonstrate that our model has achieved remarkable success in a variety of tasks related to chemical reactions. Particularly, our model has demonstrated a 25% increase in top-1 accuracy for reaction condition prediction task on the USPTO_CONDITION dataset, surpassing the strongest baseline.
2. Related Work
2.1. Molecular Representation Learning
@@
Existing methods for molecule representation learning are categorized into SMILES-based and structure-based approaches. SMILES (Weininger, 1988), as a textual representation, allows for molecule encoding with language models (Irwin et al., 2022), but it may overlook molecular topological information. Thus, there’s a growing interest in structure-based methods, which are further divided into fingerprint-based and graph neural network (GNN)-based approaches. Fingerprint-based methods, originating from the Morgan fingerprint (Morgan, 1965), face limitations due to their manual crafting and lack of end-to-end training (Ji et al., 2023), especially with complex structures and large datasets. Conversely, GNN-based learning (Jin et al., 2017; Kao et al., 2022; Ishida et al., 2021; Yang et al., 2022; Zeng et al., 2024) has gained popularity for its effectiveness.
@@
Chemical reaction representation learning, crucial for industrial applications, e.g., reactivity prediction and reaction condition optimization, has seen less focus compared to molecular representation. Current chemical reaction representation learning approaches (Lu and Zhang, 2022; Schwaller et al., 2021b) often be implemented by simple concatenation of molecular representation, or rely on the straightforward application of existing backbones without domain knowledge integration. This study presents a novel model that captures molecular differences before and after reactions and incorporates reaction center information, enhancing the robustness of chemical reaction representations.
2.2. Reaction Condition Prediction
@@
The chemical reaction condition prediction task aims to identify suitable catalysts, solvents, reagents, or other conditions for a given chemical reaction involving specific reactants and products. Existing methods can be broadly categorized into two types. The first category transforms the problem into a classification task within a predefined condition pool. GCNN (Maser et al., 2021) employs graph neural networks for multi-label classification to predict the presence of each molecule in the reaction condition combination. FPRCR (Gao et al., 2018) and Parrot (Wang et al., 2023), focusing on reaction condition combinations with fixed compositional elements, utilize fingerprinting and BERT (Devlin et al., 2019) respectively, to predict the specific reagents for each component. The second category is not constrained by a predefined reagent library. These methods (Lu and Zhang, 2022; Andronov et al., 2023) leverage language models to generate SMILES strings of the chemical reagents as reaction conditions. However, these approaches depend on manual feature selection based on expert knowledge and do not offer a generalizable prediction model with robust reaction representation capabilities.
2.3. Reaction Yield Prediction & Reaction Selectivity Prediction
@@
Reaction yield and selectivity prediction are fundamentally similar tasks, both requiring regression of a numerical value given a chemical reaction and its conditions. Consequently, many methods are applicable to both problems. Existing strategies can be divided into two primary categories: fingerprint-based and deep-learning-based. Fingerprint-based approaches construct chemical reaction representations on the basis of hand-crafted molecular fingerprints through various combinatorial strategies. DRFP (Probst et al., 2022) has designed a fingerprint that reflects the differences between reactants and products, serving as a chemical reaction representation. MFF (Sandfort et al., 2020) leverages a variety of fingerprints to enhance model performance.
@@
Deep-learning methods predominantly employ large-scale pretrained models to extract chemical reaction representations (Schwaller et al., 2021b; Lu and Zhang, 2022; Schwaller et al., 2021c; Shi et al., 2024; Han et al., 2024), aiming for enhanced generalizability. The reaction yield prediction task often grapples with noisy data, leading some studies (Chen et al., 2024b; Kwon et al., 2022) to address this by adjusting the training loss to incorporate uncertainty, thus refining model performance. In addressing reaction selectivity prediction, there is a preference for incorporating quantum chemical information. For example, the works of Li et al.; Li et al. and Zahrt et al. have incorporated descriptors such as average steric occupancy and electronic properties predicated on structures optimized via DFT calculations. Guan et al. designed a GNN that depends on the lowest-lying conformer calculated by DFT. However, the computation of these quantum chemical descriptors is exceedingly time-intensive, potentially necessitating several days for a modest sample size, which poses a challenge for its application to large-scale datasets. Furthermore, many deep-learning methods continue to directly apply backbone architectures from other domains. There is a clear demand for a backbone that can seamlessly integrate chemical information to extract potent chemical reaction representations, which remains an area for worthy exploration.
3. Preliminary
3.1. Reaction Condition Combination Generation/Prediction
In this paper, the term “reaction condition combinations” specifically refers to the combination of catalysts, solvents, and other chemical reagents. The distinction between the prediction task and the generation task lies in the fact that the prediction task utilizes a pre-defined library of chemical reagents from which appropriate combinations are selected; in contrast, the generation task does not require a pre-defined library, and the model must generate suitable molecular combinations of reagents de novo.
3.2. Reaction Selectivity Prediction
The same set of reactants can yield different products under varying conditions. The task of reaction selectivity prediction aims to forecast the proportions of different products that can be generated from a given set of reactants under specified conditions. Reaction selectivity usually includes regio-selectivity and chiral-selectivity. The former refers to the production of different products due to differences in reaction sites, while the latter refers to the production of a pair of mirror-symmetrical products. The tendency to produce a certain product is related to the corresponding intermediate state energy, which is termed . Previous work (Seeman, 1986) about transition state theory indicates that there is an exponential relationship between and chemical reaction rate, which means that the lower the energy of the intermediate state, the higher the reaction rate, and the more inclined to form the corresponding product. To be specific (Nakliang et al., 2021), the reaction ratio of producing product and could be formulated as:
where is gas constant, is the Kelvin temperature, and represents the differences in between different reaction pathways for product and . In most cases, people use the ratio of product and , i.e. , to approximate the selectivity of chemical reactions. Then the reaction selectivity prediction, which might involve multiple reactions and products, can be simplified into the prediction of of one chemical reaction with a single product.
3.3. Notations about Chemical Reactions
@@
In the realm of chemical reactions, the fundamental components are the reactants and products. These can be represented as two distinct molecular graphs, denoted by and respectively. Here, and symbolize the set of atoms, while and represent the chemical bonds that interconnect them.
@@
A cardinal principle in chemical reactions is the conservation of atoms. This principle dictates that for every atom present in the products, there exists a unique corresponding atom in the reactants. Let and with . For the sake of clarity and consistency in subsequent discussions, we stipulate that for all , atom is the unique counterpart to the atom in the products. And we define the atoms of reactants do not appear in the products as leaving group, denoted as .
@@
We further denote the set of reaction centers as , which is a subset of all the atoms from both reactants and products. An atom is considered as a reaction center as long as it meets one of the following criteria:
@@
It is an atom from the reactants (or products) whose hydrogen count is discordant with that of its corresponding atom in products (or reactants).
@@
4. Methodology
@@
We introduce a novel chemical reaction feature extractor named RAlign with an encoder-decoder architecture, as illustrated in Fig. 1. The encoder incorporates the atomic correspondence between reactants and products to generate robust features for the chemical reaction. The decoder then integrates the output features from the encoder according to the reaction center information to produce different formatted outputs tailored to downstream tasks.
4.1. Atom Aligned Encoder
@@
Understanding chemical reaction mechanisms is fundamental to developing robust representations in cheminformatics. Despite significant advancements, our grasp of these mechanisms remains incomplete, and the annotation of reaction mechanisms necessitates considerable effort from chemical experts. To navigate this challenge, we have adopted a pragmatic approach. At present, there are well-established tools (Schwaller et al., 2021a; Chen et al., 2024a) that can delineate the atomic correspondence between reactants and products in chemical reactions. Integrating this atomic mapping information into models can enhance the identification of similarities and differences between reactants and products, thereby improving the model’s capacity to understand the evolution of chemical reactions. Driven by these considerations, we introduce the Atom Aligned Encoder, a model engineered to assimilate both the chemical reaction and its associated atom-mapping data for effective encoding of chemical reactions.
@@
The Atom Aligned Encoder is structured as a series of identical blocks that iteratively refine the node and edge features, mirroring the iterative process of GNNs. Within each block, we deploy two distinct message-passing neural network (MPNN) layers for reactants and products, respectively. These layers amalgamate both node and edge features in accordance with the molecular structures, yielding intermediate node features. Subsequently, an information fusion layer is implemented to integrate intermediate features of corresponding atom pairs between reactants and products. Additionally, the intermediate node features of atoms that are absent in the products are further refined through an auxiliary feed-forward network. Ultimately, the edge features are updated contingent upon the node features of the edge termini.
@@
where represents the concatenation of features, , , and represent the initial node and edge features for reactants and products. In Eq. 3, (resp. ) represents the intermediate features of the -th block for (resp. ). Residual connections and layer normalization (Ba et al., 2016) are implemented across different layers to expedite model convergence and facilitate the stabilization of the training process. The detail implementation of initial feature extraction and the message passing network is presented in Appendix B.
4.2. Incorporating Reaction Conditions

### v2

@@ Abstract
Organic synthesis stands as a cornerstone of the chemical industry. The development of robust machine learning models to support tasks associated with organic reactions is of significant interest. However, current methods rely on hand-crafted features or direct adaptations of model architectures from other domains, which lack feasibility as data scales increase or ignore the rich chemical information inherent in reactions. To address these issues, this paper introduces RAlign, a novel chemical reaction representation learning model for various organic reaction-related tasks. By integrating atomic correspondence between reactants and products, our model discerns the molecular transformations that occur during the reaction, thereby enhancing comprehension of the reaction mechanism. We have designed an adapter structure to incorporate reaction conditions into the chemical reaction representation, allowing the model to handle various reaction conditions and to adapt to various datasets and downstream tasks. Additionally, we introduce a reaction-center-aware attention mechanism that enables the model to concentrate on key functional groups, thereby generating potent representations for chemical reactions. Our model has been evaluated on a range of downstream tasks. Experimental results indicate that our model markedly outperforms existing chemical reaction representation learning architectures on most of the datasets. We plan to open-source the code contingent upon the acceptance of the paper.
1 Introduction
Molecular design and synthesis are the core processes in the organic chemical industry, particularly in the pharmaceutical sector (Dara et al., 2022; Hasselgren & Oprea, 2024). With the surge in available data and advancements in artificial intelligence, the integration of AI into organic chemistry has achieved significant progress (Lu et al., 2018; Zhou et al., 2023). However, existing studies primarily focus on tasks related to molecular design, such as virtual screening, which typically requires modeling of individual molecules. In contrast, tasks associated with molecular synthesis like reaction conditions prediction, which necessitate modeling the various components involved in the reaction process, are still relatively unexplored and remain a challenge for humans. These molecular synthesis-related tasks rely on effective chemical reaction representations to address the relationships among molecules during complex reaction processes. Therefore, this work focuses on the backbone design for chemical reaction representation learning, with the aim of enhancing the current state of affairs.
@@ 1 Introduction
Existing chemical reaction representation learning methods can be roughly classified into two groups: fingerprint-based and deep learning-based. Fingerprint-based methods use hand-made fingerprints as molecular representations. These methods (Probst et al., 2022; Sandfort et al., 2020) employ various strategies to integrate the molecular representations of different components of a chemical reaction to derive a comprehensive chemical reaction representation. These representations are usually combined with conventional machine learning models, such as XGBoost (Chen & Guestrin, 2016), to tackle downstream tasks. Although these approaches do not require extensive computational resources and have proven effective with limited datasets, they may encounter performance limitations when scaling to larger datasets and more complex scenarios. This is attributed to the oversimplification of information inherent in the manually designed, statistically based features.
@@
Working towards more powerful chemical reaction representations, researchers have increasingly turned to deep learning techniques. Using SMILES (Weininger, 1988), chemical reactions can be encoded into a string, allowing the application of natural language processing methodologies to address the challenge of chemical reaction representation learning (Lu & Zhang, 2022; Schwaller et al., 2021b). Some studies (Maser et al., 2021; Han et al., 2024) used graph neural networks for this task, taking advantage of the natural graph structure of molecules. However, most of these methods tend to directly apply existing frameworks from other domains or perform independent feature extraction for each component of the reaction followed by a simple aggregation, potentially overlooking the rich information in complex chemical reactions.
@@
Beyond the insufficient utilization of reaction information, current methods simply regard reagents as a component of reactants for encoding when integrating reaction conditions into the reaction representation. However, this approach does not allow the model to consider any nonmolecular reaction conditions, including temperature and other environmental factors (Goodman, 2009). Furthermore, this narrow focus also impedes the applicability of current methods to datasets that provide experimental operations (e.g., stir and filter) in the form of natural language (Kearnes et al., 2021). Hence, we aim to develop a chemical reaction representation learning model that integrates a richer set of chemical information and is adaptable to various modalities of reaction conditions.
@@
To address the aforementioned shortcomings, we proposed RAlign, a powerful chemical reaction representation learning model for multiple downstream tasks. The reaction centers and the reaction process play a pivotal role in determining the outcome of the reaction (Schwaller et al., 2021b). Drawing inspiration from the imaginary transition structures of organic reactions (Fujita, 1986), we incorporate information fusion operations for corresponding atom pairs in reactants and products within our encoder. This approach explicitly models the chemical bond changes during reactions. Furthermore, we have proposed a reaction-center-aware decoder to assist the model in focusing on key functional groups. To accommodate various modalities of reaction conditions, we have employed an adapter structure to integrate these conditions into the chemical reaction representations. We evaluated our model in a variety of tasks, including reaction condition prediction, reaction yield prediction, and reaction selectivity prediction. Our model has achieved remarkable performance across all tasks, even surpassing the baselines with extensive pretraining. The contribution of this work can be summarized as:
@@
To the best of our knowledge, this work is the first to model the atomic correspondence between reactants and products in the extraction of reaction representations and introduces the first graph-based architecture specifically tailored for chemical reactions.
@@
Extensive experiments demonstrate that our model has achieved remarkable success in a variety of tasks related to chemical reactions. In particular, our model has demonstrated a 25% increase in top-1 accuracy for reaction condition prediction task on the USPTO_CONDITION dataset, surpassing the strongest baseline.
2 Related Work
2.1 Molecular Representation Learning
@@
Existing methods for molecule representation learning are categorized into SMILES-based and structure-based approaches. SMILES (Weininger, 1988), as a textual representation, allows molecule encoding with language models (Irwin et al., 2022), but it may overlook molecular topological information. Thus, there is a growing interest in structure-based methods, which are further divided into fingerprint-based and graph neural network (GNN)-based approaches. Fingerprint-based methods, originating from the Morgan fingerprint (Morgan, 1965), face limitations due to their manual crafting and lack of end-to-end training (Ji et al., 2023), especially with complex structures and large datasets. In contrast, GNN-based learning (Jin et al., 2017; Kao et al., 2022; Ishida et al., 2021; Yang et al., 2022; Zeng et al., 2024) has gained popularity due to its effectiveness.
@@
Chemical reaction representation learning, crucial for industrial applications, e.g., reactivity prediction and reaction condition optimization, has seen less focus compared to molecular representation. Current chemical reaction representation learning approaches (Lu & Zhang, 2022; Schwaller et al., 2021b) often be implemented by simple concatenation of molecular representation, or rely on the straightforward application of existing backbones without domain knowledge integration. This study presents a novel model that captures molecular differences before and after reactions and incorporates reaction center information, enhancing the robustness of chemical reaction representations.
2.2 Reaction Condition Prediction
@@
The chemical reaction condition prediction task aims to identify suitable catalysts, solvents, reagents, or other conditions for a given chemical reaction involving specific reactants and products. Existing methods can be broadly categorized into two types. The first category transforms the problem into a classification task within a predefined condition pool. GCNN (Maser et al., 2021) employs graph neural networks for multi-label classification to predict the presence of each molecule in the reaction condition combination. FPRCR (Gao et al., 2018) and Parrot (Wang et al., 2023), focusing on reaction condition combinations with fixed compositional elements, utilize fingerprinting and BERT (Devlin et al., 2019) respectively, to predict the specific reagents for each component. The second category is not constrained by a predefined reagent library. These methods (Lu & Zhang, 2022; Andronov et al., 2023) leverage language models to generate SMILES strings of chemical reagents as reaction conditions. However, these approaches are dependent on manual feature selection based on expert knowledge and do not offer a generalizable prediction model with robust reaction representation capabilities.
2.3 Reaction Yield/Selectivity Prediction
@@
Reaction yield and selectivity prediction are fundamentally similar tasks, both requiring regression of a numerical value given a chemical reaction and its conditions. Consequently, many methods are applicable to both problems. Existing strategies can be divided into two primary categories: fingerprint-based and deep learning-based. Fingerprint-based approaches construct chemical reaction representations on the basis of hand-crafted molecular fingerprints through various combinatorial strategies. DRFP (Probst et al., 2022) has designed a fingerprint that reflects the differences between reactants and products, serving as a chemical reaction representation. MFF (Sandfort et al., 2020) leverages a variety of fingerprints to enhance model performance.
@@
Deep-learning methods predominantly employ large-scale pretrained models to extract chemical reaction representations (Schwaller et al., 2021b; Lu & Zhang, 2022; Schwaller et al., 2021c; Shi et al., 2024; Han et al., 2024), with the aim of increasing generalizability. The reaction yield prediction task often grapples with noisy data, leading some studies (Chen et al., 2024a; Kwon et al., 2022) to address this by adjusting training loss to incorporate uncertainty, thus refining model performance. In addressing reaction selectivity prediction, there is a preference for incorporating quantum chemical information. For example, the works of Li et al., Li et al. and Zahrt et al. have incorporated descriptors like average steric occupancy and electronic properties predicated on structures optimized via DFT calculations. Guan et al. designed a GNN that depends on the lowest-lying conformer calculated by DFT. However, the computation of these quantum chemical descriptors is exceedingly time-intensive, potentially necessitating several days for a modest sample size, which poses a challenge for its application to large-scale datasets. Furthermore, many deep-learning methods continue to directly apply backbone architectures from other domains. There is a clear demand for a backbone that can seamlessly integrate chemical information to extract potent chemical reaction representations, which remains an area for worthy exploration.
3 Preliminary
3.1 Notations about Chemical Reactions
@@
In the realm of chemical reactions, the fundamental components are reactants and products. These can be represented as two distinct molecular graphs, denoted by and respectively. Here, and represent the set of atoms, while and represent the chemical bonds that interconnect them.
@@
A cardinal principle in chemical reactions is the conservation of atoms. This principle dictates that for every atom present in the products, there exists a unique corresponding atom in the reactants. Let and with . For clarity and consistency in subsequent discussions, we stipulate that for all , the atom is the unique counterpart to the atom in the products. And we define the atoms of reactants that do not appear in the products as the leaving group, denoted as .
@@
We further denote the set of reaction centers as , which is a subset of all atoms of both reactants and products. An atom is considered as a reaction center as long as it meets one of the following criteria:
@@
It is an atom from the reactants (or products) whose hydrogen count is discordant with that of its corresponding atom in the products (or reactants).
@@
The above definitions are extensions derived from the common definitions of reaction centers used in retrosynthesis, which are inspired by Retroformer (Wan et al., 2022).
3.2 Reaction Selectivity Prediction
The same set of reactants can produce different products under varying conditions. The task of reaction selectivity prediction aims to forecast the proportions of different products that can be generated from a given set of reactants under specified conditions. Reaction selectivity usually includes regio-selectivity and chiral-selectivity. The former refers to the production of different products due to differences in reaction sites, while the latter refers to the production of a pair of mirror-symmetrical products. The tendency to produce a certain product is related to the corresponding intermediate state energy, which is termed . Previous work (Seeman, 1986) on transition state theory indicates that there is an exponential relationship between and the chemical reaction rate, which means that the lower the energy of the intermediate state, the higher the reaction rate and the more inclined to form the corresponding product. To be specific (Nakliang et al., 2021), the reaction ratio of producing product and could be formulated as:
where is gas constant, is the Kelvin temperature, and represents the differences in between different reaction pathways for product and . In most cases, people use the ratio of the product and , i.e., , to approximate the selectivity of chemical reactions. Then the reaction selectivity prediction, which might involve multiple reactions and products, can be simplified into the prediction of of a chemical reaction with a single product.
3.3 Reaction Condition Combination Generation / Prediction
In this paper, the term “reaction condition combinations” specifically refers to the combination of catalysts, solvents, and other chemical reagents. The distinction between the prediction task and the generation task lies in the fact that the prediction task utilizes a predefined library of chemical reagents from which appropriate combinations are selected; on the contrary, the generation task does not require a predefined library and the model must generate suitable molecular combinations of reagents de novo.
4 Methodology
@@
We introduce a novel chemical reaction feature extractor named RAlign with an encoder-decoder architecture, as illustrated in Fig. 1. The encoder incorporates the atomic correspondence between reactants and products to generate robust features for the chemical reaction. The decoder then integrates the output features from the encoder according to the reaction center information to produce differently formatted outputs customized to downstream tasks.
4.1 Atom Aligned Encoder
@@
Understanding chemical reaction mechanisms is fundamental to developing robust representations in cheminformatics. Despite significant advancements, our grasp of these mechanisms remains incomplete, and the annotation of reaction mechanisms necessitates considerable effort from chemical experts. To navigate this challenge, we have adopted a pragmatic approach. Currently, there are well-established tools (Schwaller et al., 2021a; Chen et al., 2024b) that can delineate the atomic correspondence between reactants and products in chemical reactions. Integrating this atomic mapping information into models can enhance the identification of similarities and differences between reactants and products, thereby improving the model’s capacity to understand the evolution of chemical reactions. Driven by these considerations, we introduce the Atom Aligned Encoder, a model engineered to assimilate both the chemical reaction and its associated atom-mapping data for effective encoding of chemical reactions.
@@
The Atom Aligned Encoder is structured as a series of identical blocks that iteratively refine the node and edge features, mirroring the iterative process of GNNs. Within each block, we deploy two distinct message-passing neural network (MPNN) layers for reactants and products, respectively. These layers amalgamate both node and edge features in accordance with the molecular structures, yielding intermediate node features. Subsequently, an information fusion layer is implemented to integrate intermediate features of corresponding atom pairs between reactants and products. Additionally, the intermediate node features of atoms that are absent in the products are further refined through an auxiliary feedforward network. Ultimately, the edge features are updated based on the node features of the edge termini.
@@
where represents the concatenation of features, , , and represent the initial node and edge features for reactants and products. In Eq. 3, (resp. ) represents the intermediate features of the -th block for (resp. ). Residual connections and layer normalization (Ba et al., 2016) are implemented across different layers to expedite model convergence and facilitate the stabilization of the training process. The detailed implementation of initial feature extraction and the message passing network is presented in Appendix C.
4.2 Incorporating Reaction Conditions